OpenStudy (anonymous):

find the common denominator for thesse two rational expressions a^5/a^62+4a+4 and -7/a^2+7a+10

OpenStudy (daniellelovee):

is it a^62 or a^6 2+4?

OpenStudy (anonymous):

a^5 over a^2+4a+4 -7 over a^2+7a+10

OpenStudy (anonymous):

@Daniellelovee

OpenStudy (daniellelovee):

according to mathway the answer is (a+2)2(a+5)

OpenStudy (daniellelovee):

(a+2)^2(a+5)

OpenStudy (daniellelovee):

robtobey is abetter helper sorry I couldn't help so much

OpenStudy (anonymous):

\[\left\{a^2+4 a+4,a^2+7 a+10\right\} \]\[\left\{(a+2)^2,(a+2) (a+5)\right\} \]\[\left\{(a+2)^2 (a+5)\right\} \]

OpenStudy (anonymous):

x^4 over x^2-2x+4 -2 over x^2-3x+2 @robtobey

OpenStudy (anonymous):

thank you any way @daniellelove